OpenStudy (anonymous):

coefficient of the second term in the expansion of the binomial (4x+3y)^3

OpenStudy (anonymous):

(16x^2 + 24xy + 9y^2)( 4x + 3y ) =64x^3 + 48x^2y + 96x^2y + 72xy^2 + 36 xy^2 + 27y^3 =64x^3 + 144x^2y + 108xy^2 + 27y^3 If the second term is x^2y, then the answer is 144. :)

OpenStudy (welshfella):

i use the formula (r + 1)th term of (x + a)^n = nCr x^(n-r) a^r where nCr =number of combinations r in n

OpenStudy (welshfella):

= 3C1 (4x^)2* 3y = 3*16 x^2* 3y = 144 x^2 y